OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is 3x – 5 and whose width is x – 4? Answer A. 3x2 – 17x + 20 B. 3x2 + 20 C. 3x2 + 17x + 20 D. 3x2 – 20

OpenStudy (anonymous):

(3x - 5)(x-4) = 3x2 - 12x - 5x +20 = 3x2 - 17x +20 A is the answer :)

OpenStudy (anonymous):

Multiply. (x – 4)(x2 – 5x – 6)

OpenStudy (anonymous):

= x3 - 4x2 - 5x2 + 20x - 6x +24 = x3 - 9x2 + 14x + 24

OpenStudy (anonymous):

Factor completely 6x4 – 8x2 + 10x